The probability that a vehicle entering the Luray
Caverns has Canadian license plates is 0.12; the
probability that it is a camper is 0.28; and the probability
that it is a camper with Canadian license plates
is 0.09. What is the probability that
(a) a camper entering the Luray Caverns has Canadian
license plates?
(b) a vehicle with Canadian license plates entering the
Luray Caverns is a camper?
(c) a vehicle entering the Luray Caverns does not have
Canadian plates or is not a camper?

Answer by ewatrrr(24785)

p(Canadian license plates)= 0.12
p(camper)= 0.28
p( camper with Canadian license plates)= 0.09
.........
(a) P(camper with Canadian license plates) = .12
........
(b) P(a vehicle with Canadian license plates is a camper)
P(A|B) = P(A and B)/P(B)
P(Camper| Canadian Plates) = .09/.12 = .75
............
(c) P( vehicle A:does not have Canadian plates 0r B: is not a camper)
P(A 0r B) = P(A) + P(B) - P(A and B) = .88 + .72 - .81 = .80